van der waals b for oxygen is 32 cm^3/mol.compute diameter of an o2 molecule?

Grade:12th pass

1 Answers

Arun
3 years ago
b=4V 
m
​ 
 =32cm 
3
 /mol
m
​ 
 =8cm 
3
 /mol
 
For one molecule, 
V= 
6.023×10 
23
 molecules/mol
8cm 
3
 /mol
​ 
 
V=1.33×10 
−23
 cm 
3
 /molecule
But V= 
3
4
​ 
 πr 
3
 
3
4
​ 
 πr 
3
 =1.33×10 
−23
 cm 
3
 /molecule
3
 =3.17×10 
−24
 cm 
3
 /molecule
r=1.47×10 
−8
 cm
The diameter of an O 
2
​ 
  molecule =2r=2×1.47×10 
−8
 cm=2.94×10 
−8
 cm